OpenStudy (anonymous):

How do you divide this radical? 5/-3-3√3 I know that you need to rationalize the denominator. My question is: how do I multiply -3-3√3 and -3-3√3? I still end up with a radical in my denominator.

OpenStudy (anonymous):

that is because you want to multiply by the "conjugate" of the denominator, namely \[-3+\sqrt{3}\]

OpenStudy (anonymous):

you get \[(-3-\sqrt{3})(-3+\sqrt{3})=9-3=6\] and no more radical

OpenStudy (anonymous):

^We do this because it's simply reversing the "difference of two squares" case you learned in factoring.

OpenStudy (anonymous):

what yakeyglee said, although he/she is making a rather broad assumption, but that is right \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

Don't you use Multiply each number by FOIL?

OpenStudy (anonymous):

Don't you use FOIL?

OpenStudy (anonymous):

When you're multiplying the denominators?

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

foil is good for wrapping turkey leftovers, not much else a trick math teachers play on students

OpenStudy (anonymous):

it is always true that \[(a-b)(a+b)=a^2-b^2\] so \[(-3+\sqrt{3})(-3-\sqrt{3})=(-3)^2-(\sqrt{3})^2=9-3=6\]

OpenStudy (anonymous):

But -3 is not a perfect square

OpenStudy (anonymous):

true that, but i am multiplying, not factoring

OpenStudy (anonymous):

can you do that for any radical? I have another one too 2+5√3/-4+4√2

OpenStudy (anonymous):

\[\frac{2+5\sqrt{3}}{-4+4\sqrt{2}}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then multiply by \[\frac{2+5\sqrt{3}}{-4+4\sqrt{2}}\frac{(-4+4\sqrt{2})}{(-4+4\sqrt{2})}\]

OpenStudy (anonymous):

denominator will be \[(-4)^2-(4\sqrt{2})^2=16-32=-16\] the numerator will be a hot mess so i will let you do that on your own

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Hey i keep on messing up when i use your thing